OpenStudy (anonymous):

square root of 2z+9-square root of z+6=o

OpenStudy (anonymous):

\[\sqrt{2z+9 }-\sqrt{z+6}=0\]

OpenStudy (anonymous):

\[\sqrt{2z+9}=\sqrt{z+6}\]

OpenStudy (anonymous):

\[\implies \sqrt{2z+9}=\sqrt{z+6} \implies 2z+9=z+6 \implies z=-3\] Plug in the equation and check!!

OpenStudy (anonymous):

hows the work?

OpenStudy (anonymous):

Just substitute in the original equation with z=-3 to check it's the right solution!!

OpenStudy (anonymous):

It's right by the way!! So your answer is z=-3

OpenStudy (anonymous):

yea i checked!

OpenStudy (anonymous):

\[\sqrt{x+8}=x-4\]

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

x-8=x-16

OpenStudy (anonymous):

-8 -8

OpenStudy (anonymous):

=8

OpenStudy (anonymous):

Well. The first thing you should here is to get rid of the square root. To do so, you should square both sides. Can you do that?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

you should do*

OpenStudy (anonymous):

Ok. Do it and show me what you get.

OpenStudy (anonymous):

\[\sqrt({x-8})^{2}=(x-4)^{2}\]

OpenStudy (anonymous):

\[x+8=x+16\]

OpenStudy (anonymous):

are you sure about the right hand side?

OpenStudy (anonymous):

(x-4)^2 is what?

OpenStudy (anonymous):

yes because 4x4=16

OpenStudy (anonymous):

all (x-4) is raised to the power of 2. Not only the 4.. So it's like (x-4) times (x-4)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Ok.. follow my steps carefully!

OpenStudy (anonymous):

First start by squaring both sides: \[(\sqrt{x+8})^2=(x-4)^2\] you did the left hand side, it's the same as x+8.. the right hand side is: \[(x-4)^2=x^2-8x+16\]

OpenStudy (anonymous):

Now, the equation will look like: \[x+8=x^2-8x+16 \implies x^2-9x+8=0\] Do you know how to solve a quadratic equation?

OpenStudy (anonymous):

Just factorize the expression x^2-9x+8 to get: \[x^2-9x+8=0 \implies (x-8)(x-1)=0 \implies x=1, x=8\] We have two values for x. Check for each one!

OpenStudy (anonymous):

You will find that only x=8 is a solution of the equation.